ABC PLC, which has a year-end of 31st December 2012, issued 2 for 3 bonus shares on 30th June 2012. Net income on 31 December 2011 is $7 000 000 and on 31 December 2012 is $8 000 000 and there were 3 000 000 shares on 1 January 2012. What is the earnings per share for 2012 taking into consideration the bonus issue?

$1.33

$1.40

$1.60
Answer #1

Shares end of year = shares beginning of year(1+bonus ratio)

=3000000*(1+2/3)

=

5000000

EPS = net income/number of shares = 8000000/5000000 = 1.6